# A solution to the Cauchy dual subnormality problem for a cyclic analytic $2$-isometry with defect operator of rank two

### v1

@@ 2. Case (ii)
      There exist $x_{0}\in(\beta,5)$ such that $g(x_{0})>0$.
      But, $g(5)<0$ and hence there exist $\gamma\in(x_{0},5)$ such that $g(\gamma)=0$. Thus, we get, $g(\frac{1}{\gamma})=0$. This implies that $g$ has more than $4$ roots, which is a contradiction.
    Thus, in any case, we get a contradiction. Hence, $g$ has no root in $(1,5)$. Therefore, $b>5$.
    Hence, if $\cos\theta<0.6$ then we get $D<0$. But this is a contradiction, since we know that $s,t$ are the roots of the polynomial $h(x)$. This is a contradiction, and hence, we must have $\arg(\alpha_{1})\neq\arg(\alpha_{2}).$
@@
    Let $\cos\theta>0.6.$ We need to check whether there exist $s,t\in\mathbb{R}$ such that $\alpha_{1}=s(1+\xi)$ and $\alpha_{2}=t(1+\xi)$.
    In this case, we get $\alpha_{1}+\alpha_{2}=(s+t)(1+\xi)=a(1+\xi)$ and $\alpha_{1}\alpha_{2}=st(1+\xi)^{2}=b\xi$.
    Thus we get $s+t=a$ and $st=\displaystyle\frac{b\xi}{(1+\xi)^{2}}=\frac{b}{|1+\xi|^{2}}=\frac{b}{2(1+\cos\theta)}$.
    Consider a polynomial h(x)=(xs)(xt)h(x)=(x-s)(x-t).
    The discriminant $D$ of polynomial $h(x)$ becomes
    $D=\displaystyle\frac{9b^{2}}{(b+1)^{2}}-\frac{2b}{1+\cos\theta}$
    We have, $\cos\theta>0.6$, which gives $\frac{-1}{1+\cos\theta}>\frac{-1}{1.6}$.
    Thus,
    $D$ $=\displaystyle\frac{9b^{2}}{(b+1)^{2}}-\frac{2b}{1+\cos\theta}$
    $>\displaystyle\frac{9b^{2}}{(b+1)^{2}}-\frac{2b}{1.6}=\displaystyle\frac{9b^{2}}{(b+1)^{2}}-\frac{5b}{4}$
    $=\displaystyle\frac{36b^{2}-5b(b^{2}+2b+1)}{4(b+1)^{2}}=\displaystyle\frac{-b(5b^{2}-26b+5)}{4(b+1)^{2}}$
    $=\displaystyle\frac{-b(5b-1)(b-5)}{4(b+1)^{2}}$
    Thus, $D>\displaystyle\frac{-b(5b-1)(b-5)}{4(b+1)^{2}}$
@@
    Now, for $\cos\theta>0.6$, from equation 3.13, we get
    $g(1)>0,g(5)>0,g(6)>0$
    Note that $g(4)<0$ if $\cos\theta<\displaystyle\frac{87}{88}$ and
    $g(3)<0$ if $\cos\theta<\displaystyle\frac{62}{66}$.
    Thus, if $0.6<\cos\theta<\frac{62}{66}$ then $g(3)<0$ which implies that there exist $\beta_{0}\in(1,3)$ and $\gamma_{0}\in(3,5)$ such that $g(\beta_{0})=0=g(\gamma_{0})$. Also, by property of $g$, we get, $g(\frac{1}{\beta_{0}})=0=g(\frac{1}{\gamma_{0}}$.
    This implies that all the roots of $g$ are less than $5$.
    Now, if $\frac{62}{66}<\cos\theta<\frac{87}{88}$ then $g(4)<0$. A similar argument as above shows that all the roots of $g$ are less than $5$.
    Finally, if $\frac{87}{88}<\cos\theta<1$ then note that $g(3)>0$ as well as $g(4)>0$.
@@
    Thus, $g(2+\sqrt{3})<0$ since $\cos\theta<1$.
    This assures that $g$ has two roots in the interval $(3,4)$ and hence another two roots in $(0,1)$.
    Thus, all the roots of $g$ are less than $5$.
    Hence, in any case, we have $b<5$.
    Hence, $D>0$.
    This implies that the polynomial $h(x)$ has two distinct real roots $s,t$.
    Hence, there exist $s,t\in\mathbb{R}$ such that $\alpha_{1}=s(1+\xi)$ and $\alpha_{2}=t(1+\xi)$.
    Therefore, $\arg(\alpha_{1})=\arg(\alpha_{2})$.
    Further, note that since $s,t$ are distinct real roots, we must have $|\alpha_{1}|\neq|\alpha_{2}|$. Thus, by suitably choosing notations for $\alpha_{1}$ and $\alpha_{2}$ we get $k>1$ such that $\alpha_{2}=k\alpha_{1}$.
@@
    It can be verified that if $\cos\theta=0.6$ then $f(2+i)=0$. In fact, we get,
    $f(z)=(z-(2+i))^{2}\left(z-\left(\frac{1}{2-i}\right)\right)^{2}.$
    Therefore, $\alpha_{1}=\alpha_{2}=2+i$.
We first prove the main theorem in the case where $\cos\theta<0.6$. The proof relies on the technique discussed in [16]. The computational details in this case have been discussed below.
The polynomial $q(z)$ obtained in the equation (3.9) allows us to compute $O_{\mu}(z)$ as follows:
$O_{\mu}(z)=\displaystyle\frac{p(z)}{q(z)}$
where $p(z)=\displaystyle\frac{e^{i\theta_{0}}}{\sqrt{d}}(z-1)(z-\xi)$ and $q(z)=(z-\alpha_{1})(z-\alpha_{2})$.
$O_{\mu}(0)=\displaystyle\frac{p(0)}{q(0)}=\frac{e^{i\theta_{0}}(-1)(-\xi)}{\sqrt{d}\alpha_{1}\alpha_{2}}=\frac{e^{i\theta_{0}}}{\sqrt{d}~b}$.
@@
$p(z)=\displaystyle\frac{1}{\sqrt{d}}(z-1)(z-\xi)\text{ and }O_{\mu}(z)=\displaystyle\frac{1}{\sqrt{d}}\frac{(z-1)(z-\xi)}{q(z)}.$ (3.14)
A close look at [16, Theorem 6.4] suggest that we need to first compute the matrix $B=\left(\left[\langle f_{i},f_{j}\rangle_{\mu}\right]_{1\leq i,j\leq 2}\right)^{-1}$ present in the equation (2.6). We thus proceed for the computations of this matrix through the following lemma.
@@
Let matrix $C=\left[c_{ij}\right]_{1\leq i,j\leq 2}$, where $c_{ij}=\langle f_{i},f_{j}\rangle_{\mu}$, $1\leq i,j\leq 2$ and $f_{i}$ as defined in (2.5). Then
@@
    Matrix $B$ is given by $B=\begin{pmatrix}a-1&-\displaystyle\frac{c_{12}}{\Delta}\\ -\displaystyle\frac{\overline{c_{12}}}{\Delta}&a-1\end{pmatrix}$, where $c_{12}=\displaystyle\frac{dq(1)^{2}}{(\xi-1)^{3}}$.
@@
    Thus, $O_{\mu}^{\prime}(1)=\displaystyle\frac{1}{\sqrt{d}}\frac{1-\xi}{q(1)}$ and $O_{\mu}^{\prime}(\xi)=\displaystyle\frac{1}{\sqrt{d}}\frac{\xi-1}{q(\xi)}$
@@
    This gives the following relation : O_μ’(ξ)=ξ O_μ’(1).
@@
    $f_{1}(z)=\displaystyle\frac{q(1)}{1-\xi}\frac{z-\xi}{q(z)}\text{ and }f_{2}(z)=\displaystyle\frac{z-1}{q(z)}\frac{q(\xi)}{\xi-1}.$ (3.15)
@@
    Hence, by equation (2.7) we get, c_11=f_1’(1)=11-ξ-q’(1)q(1). But, $q(z)=z^{2}-a(1+\xi)z+b\xi$, hence, we get,
    $c_{11}$ $=\displaystyle\frac{1}{1-\xi}+\frac{a-2+a\xi}{1-a+(b-a)\xi}$
    $=\Re\left(\displaystyle\frac{1}{1-\xi}\right)+\frac{(a-2)(1-a)+a(b-a)+a(1-a)\cos\theta+(a-2)(b-a)\cos\theta}{x^{2}+y^{2}+2xy\cos\theta},$
    where x=1ax=1-a and y=bay=b-a. Thus we get,
    $x^{2}+y^{2}+2xy\cos\theta=\displaystyle\frac{g(b)+2(1-\cos\theta)b(b^{2}+2b+1)}{(b+1)^{2}}=2b(1-\cos\theta)$
    Hence,
    $c_{11}=\displaystyle\frac{1}{2}+\frac{(a-2)(1-a)+a(b-a)+a(1-a)\cos\theta+(a-2)(b-a)\cos\theta}{2b(1-\cos\theta)}$
    As the numerator of $c_{11}$ has (a1)(a-1) as a factor, this along with the observation that b(3a)=ab(3-a)=a gives,
    $c_{11}=(a-1)\left[(a-3)\displaystyle\frac{1+\cos\theta}{1-\cos\theta}+\frac{3}{a(1-\cos\theta)}\right]$ (3.16)
    We once again use equations (2.7) and (2.5) to get, c22=ξf2’(ξ) where f2(z)=(z-1)dq(z)Oμ’(ξ). We further have, $\frac{O_{\mu}(z)}{z-\xi}=\frac{1}{\sqrt{d}}\frac{z-1}{q(z)}.$
    $f_{2}^{\prime}(z)=\displaystyle\frac{1}{\sqrt{d}O_{\mu}^{\prime}(\xi)}\left[\frac{q(z)-(z-1)q^{\prime}(z)}{q(z)^{2}}\right]$
    Hence, $f_{2}^{\prime}(\xi)=\displaystyle\frac{1}{\xi-1}-\frac{q^{\prime}(\xi)}{q(\xi)}$
    This gives, $c_{22}=\xi f_{2}^{\prime}(\xi)=\displaystyle\frac{\xi}{\xi-1}-\frac{\xi q^{\prime}(\xi)}{q(\xi)}$
    Note that $\frac{\xi}{\xi-1}=\frac{1}{1-\overline{\xi}}$, $\overline{\xi}q^{\prime}(\xi)=\overline{q^{\prime}(1)}$.
    Thus, we get, c22=11-ξ-ξq’(ξ)ξ2q(1)=11-ξ-q’(1)q(1)=c11.
    Now we use equation (2.8) to obtain the following: c12=1Oμ’(1)Oμ’(ξ)(1-ξ) =d q(1)2(ξ-1)3
    Hence, $|c_{12}|^{2}=\displaystyle\frac{1}{2(1-\cos\theta)}.$
    Observe that $c_{21}=\overline{c_{12}}$
    We have, $\Delta=\det(C)=c_{11}c_{22}-|c_{12}|^{2}$
    Hence,
@@
    We prove that [ (a-1)2(a(a-3)(1+cosθ)+3)a(1-cosθ) -a2(a(a-3)(1+cosθ)+3)]=1. It is enough to prove that
@@
    But, we know that $a-1=\displaystyle\frac{2b-1}{b+1}$, and thus the above condition reduces to
    $9(3b^{2}\cos\theta+5b^{2}-6b\cos\theta-2b+2)g(b)=0.$
    This is true since $b$ is a root of the polynomial $g(z).$
@@
    $\Delta=\det(C)=\left[(a-3)\displaystyle\frac{1+\cos\theta}{1-\cos\theta}+\frac{3}{a(1-\cos\theta)}\right].$ (3.17)
@@
    The equations (3.16) and (3.17) implies $b_{11}=c_{11}/\Delta=a-1$.
@@
    $B=\begin{pmatrix}a-1&-\displaystyle\frac{c_{12}}{\Delta}\\ -\displaystyle\frac{\overline{c_{12}}}{\Delta}&a-1\end{pmatrix}$, (3.18)
@@
The following theorem is a generalization of [30, Theorem 8], developed in the case of $\{1,i\}$. We use proof of Theorem 3.1 to find $B=(b_{1},b_{2})$ such that $D(\mu)$ coincides with $H(B)$ with equality of norms.
@@
Let $\mu$ be a positive, Borel measure on the unit circle $\mathbb{T}$ of the form $\mu=\delta_{1}+\delta_{\xi}$. Then the Dirichlet space $D(\mu)$ coincides with the de Branges-Rovnyak space $H(B)$ with $B=(b_{1},b_{2})$ and $b_{j}=\displaystyle\frac{p_{j}}{q}$, $j=1,2$ where
@@
Let $\xi_{1}=1$, $\xi_{2}=\xi$, $B=(b_{ij})_{1\leq i,j\leq 2}$ be the matrix obtained in Lemma 3.6. Also, let $O_{\mu}(z)$ as defined by equation (3.14).
Since the multiplication operator $M_{z}$ on $D(\mu)$ is an analytic norm increasing operator (see [32, Theorem 3.6]), and the reproducing kernel for $D(\mu)$ is normalized. By [16, Lemma 3.4], there exists a positive semi-definite kernel $\eta:\mathbb{D}\times\mathbb{D}\to\mathbb{C}$ such that
$\eta(z,0)=0,~~~k(z,w)=\displaystyle\frac{1-\eta(z,w)}{1-z\overline{w}},~~z,w\in\mathbb{D}$
We rewrite the expression for $K(z,w)$ using equations (2.3), (2.4) and Theorem 2.1 as
@@
$=\displaystyle\frac{1}{1-z\overline{w}}\frac{p(z)\overline{p(w)}}{q(z)\overline{q(w)}}+\sum_{i,j=1}^{2}\overline{b_{ji}}f_{j}(z)\overline{f_{i}(w)}$
@@
$q(z)\eta(z,w)\overline{q(w)}=q(z)\overline{q(w)}-p(z)\overline{p(w)}\left(1+(1-z\overline{w})\displaystyle\sum_{i,j=1}^{2}\frac{\overline{b_{ji}}}{O_{\mu}^{\prime}(\xi_{j})\overline{O_{\mu}^{\prime}(\xi_{i})}}\frac{1}{(z-\xi_{j})(\overline{w}-\overline{\xi_{i}})}\right)$
Since, the right hand side is a polynomial in $z$ and $\overline{w}$, there exists a matrix $\hat{A}=(a_{ij})_{0\leq i,j\leq 2}$ such that the above expression is equal to $\sum_{i,j=0}^{2}a_{ij}z^{i}\overline{w}^{j}$, $z,w\in\mathbb{D}$.
As this expression at $w=0$ is 0 for each $z\in\mathbb{D}$, we obtain a matrix $A$ from $\hat{A}$ by deleting first row and first column of matrix $\hat{A}$.
We present here the computations to obtain matrix $A$.
We begin with simplification of the right hand side.
@@
$\hskip 65.44142pt+\frac{\overline{b_{12}}(z\overline{w}-z-\xi\overline{w}+\xi)}{\xi O_{\mu}^{\prime}(1)^{2}}+\frac{\overline{b_{21}}(z\overline{w}-z\overline{\xi}-\overline{w}+\overline{\xi})}{\xi O_{\mu}^{\prime}(\xi)^{2}}\Bigg]$
Now, $a_{11}=\text{coefficient of }z\overline{w}$
Hence,
$a_{11}=(a^{2}-b)(|1+\xi|^{2})+\displaystyle\frac{2b}{\Delta}$
$a_{22}=\text{coefficient of }z^{2}\overline{w}^{2}$ and hence,
$a_{22}=\displaystyle 1-3b+2ab-\frac{b}{\Delta}.$
$a_{12}=\text{coefficient of }z\overline{w}^{2}$ and hence,
$a_{12}=-b(1+\xi).$
Hence, we have matrix $A=(a_{ij})_{1\leq i,j\leq 2}$ which is positive definite and hence by applying Cholesky’s decomposition we get an upper triangular matrix $P$ such that $A=P^{*}P$, where
@@
$\displaystyle p_{1}(z)$ $=\langle P(z,z^{2})^{T},(1,0)\rangle=\sqrt{a_{11}}~z+\frac{a_{12}}{\sqrt{a_{11}}}z^{2}$ (3.19)
$\displaystyle p_{2}(z)$ $=\langle P(z,z^{2})^{T},(0,1)\rangle=p_{22}z^{2}=\sqrt{\frac{a_{11}a_{22}-|a_{12}|^{2}}{a_{11}}}~z^{2}.$ (3.20)
Therefore, we conclude that the Dirichlet type space $D(\mu)$ coincides with de Branges-Rovnyak space $H(B)$ with $B=(b_{1},b_{2})$ and $b_{j}=\displaystyle\frac{p_{j}}{q}$, $j=1,2.$

### v2

@@ 2. Case (ii)
      Suppose there exists $x_{0}\in(\beta,5)$ such that $g(x_{0})>0$.
      But, $g(5)<0$ and hence there exists $\gamma\in(x_{0},5)$ such that $g(\gamma)=0$. Thus, we get, $g(\frac{1}{\gamma})=0$. This implies that $g$ has more than $4$ roots, which is a contradiction.
    Hence, $g$ has no root in the interval $(1,5)$. Therefore, $b>5$.
    Thus, if $\cos\theta<0.6$ then $D<0$ contradicting the assumption that $s,t$ are the real roots of the polynomial $h(x)$. Hence,
    $\arg(\alpha_{1})\neq\arg(\alpha_{2}).$
@@
    Let $\cos\theta>0.6.$ We need to check whether there exist $s,t\in\mathbb{R}$ such that $\alpha_{1}=s(1+\xi)$ and $\alpha_{2}=t(1+\xi)$. Indeed, it is enough to show that there exist $s,t\in\mathbb{R}$ such that $\alpha_{1}+\alpha_{2}=(s+t)(1+\xi)=a(1+\xi)$ and $\alpha_{1}\alpha_{2}=st(1+\xi)^{2}=b\xi$.
    Following the steps in (1), we have the discriminant $D$ of polynomial $h(x)$ as
    $D=\displaystyle\frac{9b^{2}}{(b+1)^{2}}-\frac{2b}{1+\cos\theta}.$
    Here, we have, $\cos\theta>0.6$, which gives $\frac{-1}{1+\cos\theta}>\frac{-1}{1.6}$.
    Thus, $D>\displaystyle\frac{-b(5b-1)(b-5)}{4(b+1)^{2}}.$
@@
    Now, for $\cos\theta>0.6$, from equation (3.13), we get
    $g(1)>0,g(5)>0,g(6)>0.$
@@
    Thus, $g(2+\sqrt{3})<0$ since $\cos\theta<1$. This assures that $g$ has two roots in the interval $(1,5)$ and hence another two roots in the interval $(0,1)$. Thus, all the roots of $g$ are less than $5$. Hence, we have $b<5$ implying $D>0$. Hence, there exist distinct real numbers $s,t$ such that $\alpha_{1}=s(1+\xi)$ and $\alpha_{2}=t(1+\xi)$. Therefore, $\arg(\alpha_{1})=\arg(\alpha_{2})$.
    Further, note that since $s,t$ are distinct real roots, we must have $|\alpha_{1}|\neq|\alpha_{2}|$.
@@
    It can be verified that if $\cos\theta=0.6$ then $2+i$ is a repeated root of $f(z)$. Therefore, $\alpha_{1}=\alpha_{2}=2+i$.
The proof of the main theorem relies on the technique discussed in [13]. The computational details in this case have been discussed below.
The polynomial $q(z)$ obtained in the equation (3.9) allows us to compute $O_{\mu}(z)$ as follows: $O_{\mu}(z)=\displaystyle\frac{p(z)}{q(z)}$ where $p(z)=\displaystyle\frac{e^{i\theta_{0}}}{\sqrt{d}}(z-1)(z-\xi)$ and
$q(z)=(z-\alpha_{1})(z-\alpha_{2})$.
Now $O_{\mu}(0)=\displaystyle\frac{p(0)}{q(0)}=\frac{e^{i\theta_{0}}(-1)(-\xi)}{\sqrt{d}\alpha_{1}\alpha_{2}}=\frac{e^{i\theta_{0}}}{\sqrt{d}~b}$.
@@
(3.14) $p(z)=\displaystyle\frac{1}{\sqrt{d}}(z-1)(z-\xi)\text{ and }O_{\mu}(z)=\displaystyle\frac{1}{\sqrt{d}}\frac{(z-1)(z-\xi)}{q(z)}.$
A close look at [13, Theorem 6.4] suggests that we need to first compute the matrix $\left(\left[\langle f_{i},f_{j}\rangle_{\mu}\right]_{1\leq i,j\leq 2}\right)^{-1}$ (say $B$) in the equation (2.6). We thus proceed for the computations of the entries of this matrix through the following lemma.
@@
Let matrix $C=\left[c_{ij}\right]_{1\leq i,j\leq 2}$, where $c_{ij}=\langle f_{i},f_{j}\rangle_{\mu}$, $1\leq i,j\leq 2$ and $f_{1},f_{2}$ as defined in (2.5). Then
@@
    Matrix $B$ is given by $B=\begin{pmatrix}a-1&-\displaystyle\frac{c_{12}}{\Delta}\\ -\displaystyle\frac{\overline{c_{12}}}{\Delta}&a-1\end{pmatrix}$, where $c_{12}$ is given by $c_{12}=\displaystyle\frac{dq(1)^{2}}{(\xi-1)^{3}}$.
@@
    Thus, $O_{\mu}^{\prime}(1)=\displaystyle\frac{1}{\sqrt{d}}\frac{1-\xi}{q(1)}$ and $O_{\mu}^{\prime}(\xi)=\displaystyle\frac{1}{\sqrt{d}}\frac{\xi-1}{q(\xi)}.$
@@
    This gives the relation : $O_{\mu}^{\prime}(\xi)=\overline{\xi}~\overline{O_{\mu}^{\prime}(1)}.$
@@
    (3.15) $f_{1}(z)=\displaystyle\frac{q(1)}{1-\xi}\frac{z-\xi}{q(z)}\text{ and }f_{2}(z)=\displaystyle\frac{z-1}{q(z)}\frac{q(\xi)}{\xi-1}.$
@@
    Hence, by equation (2.7) we get, $c_{11}=f_{1}^{\prime}(1)=\displaystyle\frac{1}{1-\xi}-\frac{q^{\prime}(1)}{q(1)}.$
    But, $q(z)=z^{2}-a(1+\xi)z+b\xi$, hence, we get,
    $c_{11}=\displaystyle\frac{1}{2}+\frac{(a-2)(1-a)+a(b-a)+a(1-a)\cos\theta+(a-2)(b-a)\cos\theta}{2b(1-\cos\theta)}.$
    It can be seen that the numerator of $c_{11}$ has (a1)(a-1) as a factor, this along with the observation that b(3a)=ab(3-a)=a gives,
    (3.16) $c_{11}=(a-1)\left[(a-3)\displaystyle\frac{1+\cos\theta}{1-\cos\theta}+\frac{3}{a(1-\cos\theta)}\right].$
    We once again use equations (2.7) and (2.5) to get, c_22=ξf_2’(ξ) where f_2(z)=(z-1)dq(z)Oμ’(ξ). It is easy to see that $f_{2}^{\prime}(z)=\displaystyle\frac{1}{\sqrt{d}O_{\mu}^{\prime}(\xi)}\left[\frac{q(z)-(z-1)q^{\prime}(z)}{q(z)^{2}}\right].$
    Hence, $f_{2}^{\prime}(\xi)=\displaystyle\frac{1}{\xi-1}-\frac{q^{\prime}(\xi)}{q(\xi)}.$ Thus, $c_{22}=\xi f_{2}^{\prime}(\xi)=\displaystyle\frac{\xi}{\xi-1}-\frac{\xi q^{\prime}(\xi)}{q(\xi)}.$ It can be seen that c22=11-ξ-ξq’(ξ)ξ2q(1)=11-ξ-q’(1)q(1)=c11.
    Now we use equation (2.8) to obtain the following: c12=1Oμ’(1)Oμ’(ξ)(1-ξ) =d q(1)2(ξ-1)3. Thus, after simplification we get, $|c_{12}|^{2}=\displaystyle\frac{1}{2(1-\cos\theta)}.$
    Observe that $c_{21}=\overline{c_{12}}.$
    We express $\Delta=\det(C)=c_{11}c_{22}-|c_{12}|^{2}$ as follows:
@@
    We now prove that [ (a-1)2(a(a-3)(1+cosθ)+3)a(1-cosθ) -a2(a(a-3)(1+cosθ)+3)]=1. It is enough to prove that
@@
    But, we know that $a-1=\frac{2b-1}{b+1}$, and thus the above equation reduces to $9(3b^{2}\cos\theta+5b^{2}-6b\cos\theta-2b+2)g(b)=0.$
    This follows from the fact that $b$ is a root of the polynomial $g(z).$
@@
    (3.17) $\Delta=\det(C)=\left[(a-3)\displaystyle\frac{1+\cos\theta}{1-\cos\theta}+\frac{3}{a(1-\cos\theta)}\right].$
@@
    The equations (3.16) and (3.17) imply that $b_{11}=c_{11}/\Delta=a-1$.
@@
    (3.18) $B=\begin{pmatrix}a-1&-\displaystyle\frac{c_{12}}{\Delta}\\ -\displaystyle\frac{\overline{c_{12}}}{\Delta}&a-1\end{pmatrix}$,
@@
We now derive the following observations about $b$, $a$ and $\Delta$ as functions of $\theta$ when $\cos\theta<0.6.$ These observations are used in the proof of the main theorem.
Lemma 3.7.
Let $\xi=\cos\theta+i\sin\theta\in\mathbb{T}$ $(\xi\neq\pm 1)$. Let $\alpha_{1},\alpha_{2}$ be two complex numbers outside closed unit disc, as obtained in part (3) of Lemma 3.4. Let $b=|\alpha_{1}\alpha_{2}|$, $a=\frac{3b}{b+1}$ and $\Delta$ as obtained in part (1) of Lemma 3.6. Let $t_{0}=\cos^{-1}(0.6)$. Then, for $\theta\in[t_{0},\pi/2],$
  1. (1)
    $b$ is an increasing function of $\theta.$
  2. (2)
    $a$ is an increasing function of $\theta.$
  3. (3)
    $0.7<\Delta<1.10056.$
Proof.
  1. (1)
    We note from part (7) of Lemma 3.4 that $b$ is a root of
    $g(x)=x^{4}-Ax^{3}+Bx^{2}-Ax+1,$ where $A=6+2\cos\theta,B=4+14\cos\theta.$
    Thus, $\frac{g(x)}{x^{2}}=h(y),$ where $h(y)=y^{2}-Ay+B-2,$ where $y=x+\frac{1}{x}.$
    $h(y)=0\implies y=\frac{A\pm\sqrt{A^{2}-4B+8}}{2}.$
    It can be seen that $\frac{A+\sqrt{A^{2}-4B+8}}{2}>2.$ Let $y_{1}=\frac{A+\sqrt{A^{2}-4B+8}}{2}>2.$
    Let $c(\theta)=\cos\theta$ be a function defined on $t_{0},\pi/2].$
    Then we get,
    $y_{1}(c)=3+c+\sqrt{c^{2}-8c+7}$ and hence $\frac{dy_{1}}{dc}=1+\frac{c-4}{\sqrt{c^{2}-8c+7}}.$
    Now, $|c-4|>\sqrt{c^{2}-8c+7}$ and $c<4$ gives, $\frac{dy_{1}}{dc}<0$ and hence $y_{1}$ is a decreasing function of $c$. Further, $c$ is a decreasing function of $\theta$ and hence, $y_{1}$ is an increasing function of $\theta.$ This proves that $\frac{y_{1}(\theta)+\sqrt{y_{1}(\theta)^{2}-4}}{2}$ is an increasing function of $\theta.$
    We know that $b$ is a root of $y=x+\frac{1}{x}.$
    Thus we have $b(\theta)=\frac{y_{1}(\theta)\pm\sqrt{y_{1}(\theta)^{2}-4}}{2}.$
    As $b>5,$ we must have $b(\theta)=\frac{y_{1}(\theta)+\sqrt{y_{1}(\theta)^{2}-4}}{2}.$ Hence, $b$ is an increasing function of $\theta.$
  2. (2)
    We consider $a$ as a function of $b$ given by $a(b)=\frac{3b}{b+1}$. It can be seen that $a^{\prime}(b)>0$ and hence $a$ is an increasing function of $b$. Thus, $a$ is an increasing function of $\theta$ on $t_{0},\pi/2].$
  3. (3)
    We know that for $\theta=t_{0},$ $b=5$ (See Part (3) of Lemma 3.5) and for $\theta=\frac{\pi}{2},$ $b=5.46269$ (Refer [28, Appendix A.2]). Since, $a$ is an increasing function of $b$, we get, $a\in[2.5,2.5357]$ whenever
    $\theta\in[t_{0},\pi/2].$
    We have $\Delta=(a-3)\displaystyle\frac{1+\cos\theta}{1-\cos\theta}+\frac{3}{a(1-\cos\theta)}.$
    For fixed $a\in[2.5,2.5357]$, we have, $\Delta^{\prime}(\theta)=\frac{-(2a^{2}-6a+3)\sin\theta}{a(1-\cos\theta)^{2}}<0.$
    Thus, for fixed $a$, $\Delta$ is a decreasing function of $\theta.$
    For fixed $\theta\in[t_{0},\pi/2]$, we have, $\Delta^{\prime}(a)=\frac{a^{2}(1+\cos\theta)-3}{a^{2}(1-\cos\theta)}>0.$
    Thus, for fixed $\theta$, $\Delta$ is an increasing function of $a.$ Now, for any $\theta\in[t_{0},\pi/2]$ and $a\in[2.5,2.5357]$, we get,
    $\Delta(\theta,a)>\Delta(\theta,2.5)>\Delta(\pi/2,2.5)=0.7.$ and
    $\Delta(\theta,a)<\Delta(\theta,2.5357)<\Delta(t_{0},2.5357)=1.10056.$
    Thus, we get, $0.7<\Delta<1.10056.$
The following theorem is a generalization of [28, Theorem 8], developed in the case of $\{1,i\}$. We use proof of Theorem 3.1 to find $B=(b_{1},b_{2})$ such that $D(\mu)$ coincides with $H(B)$ with equality of norms.
@@
Let $\xi=\cos\theta+i\sin\theta\in\mathbb{T}$ $(\xi\neq\pm 1)$. Let $\mu$ be a finite, positive, Borel measure on the unit circle $\mathbb{T}$ of the form $\mu=\delta_{1}+\delta_{\xi}$. Then the Dirichlet-type space $D(\mu)$ coincides with the de Branges-Rovnyak space $H(B)$ with equality of norms where $B=(b_{1},b_{2})$ and $b_{j}=\displaystyle\frac{p_{j}}{q}$, $j=1,2$ where
@@
Let $\xi_{1}=1$, $\xi_{2}=\xi$, $B=(b_{ij})_{1\leq i,j\leq 2}$ be the matrix obtained in Lemma 3.6. Also, let $O_{\mu}(z)$ be defined by equation (3.14).
The operator $M_{z}$ on $D(\mu)$ is analytic, norm increasing (see [30, Theorem 3.6]), and the reproducing kernel for $D(\mu)$ is normalized. Therefore, by [13, Lemma 3.4], there exists a positive semi-definite kernel $\eta:\mathbb{D}\times\mathbb{D}\to\mathbb{C}$ such that
$\eta(z,0)=0,~~~k(z,w)=\displaystyle\frac{1-\eta(z,w)}{1-z\overline{w}},~~z,w\in\mathbb{D}.$
We rewrite the expression for $K(z,w)$ using equations (2.3), (2.4) and
Theorem 2.1 as
@@
$=\displaystyle\frac{1}{1-z\overline{w}}\frac{p(z)\overline{p(w)}}{q(z)\overline{q(w)}}+\sum_{i,j=1}^{2}\overline{b_{ji}}f_{j}(z)\overline{f_{i}(w)}.$
@@
$q(z)\eta(z,w)\overline{q(w)}$
$q(z)\overline{q(w)}-p(z)\overline{p(w)}\left(1+(1-z\overline{w})\displaystyle\sum_{i,j=1}^{2}\frac{\overline{b_{ji}}}{O_{\mu}^{\prime}(\xi_{j})\overline{O_{\mu}^{\prime}(\xi_{i})}}\frac{1}{(z-\xi_{j})(\overline{w}-\overline{\xi_{i}})}\right).$
Since, the right hand side is a polynomial in $z$ and $\overline{w}$, there exists a matrix $\hat{A}=(a_{ij})_{0\leq i,j\leq 2}$ such that the above expression is equal to $\sum_{i,j=0}^{2}a_{ij}z^{i}\overline{w}^{j}$, $z,w\in\mathbb{D}$. The matrix $\hat{A}$ is positive semi-definite.
As this expression at $w=0$ is 0 for each $z\in\mathbb{D}$, we obtain a matrix $A$ from $\hat{A}$ by deleting first row and first column of matrix $\hat{A}$. Therefore, matrix $A$ is also positive semi-definite.
We present here the computations to obtain matrix $A$. We begin with simplification of the right hand side.
@@
$\hskip 65.44142pt+\frac{\overline{b_{12}}(z\overline{w}-z-\xi\overline{w}+\xi)}{\xi O_{\mu}^{\prime}(1)^{2}}+\frac{\overline{b_{21}}(z\overline{w}-z\overline{\xi}-\overline{w}+\overline{\xi})}{\xi O_{\mu}^{\prime}(\xi)^{2}}\Bigg].$
We get, $a_{11}=\text{coefficient of }z\overline{w}=$ $(a^{2}-b)(|1+\xi|^{2})+\displaystyle\frac{2b}{\Delta}.$
$a_{22}=\text{coefficient of }z^{2}\overline{w}^{2}$ $1-3b+2ab-\frac{b}{\Delta}.$
$a_{12}=\text{coefficient of }z\overline{w}^{2}=-b(1+\xi).$
Note that $a_{11}$ and $det(A)=a_{11}a_{22}-|a_{12}|^{2}$ are non-negative real numbers. Thus, $a_{22}$ is also a non-negative real number.
As matrix $A=(a_{ij})_{1\leq i,j\leq 2}$ is positive semi-definite, by applying Cholesky’s decomposition we get an upper triangular matrix $P$ such that $A=P^{*}P$, where
@@
(3.19) $\displaystyle p_{1}(z)$ $=\langle P(z,z^{2})^{T},(1,0)\rangle=\sqrt{a_{11}}~z+\frac{a_{12}}{\sqrt{a_{11}}}z^{2},$
(3.20) $\displaystyle p_{2}(z)$ $=\langle P(z,z^{2})^{T},(0,1)\rangle=p_{22}z^{2}=\sqrt{\frac{a_{11}a_{22}-|a_{12}|^{2}}{a_{11}}}~z^{2}.$
Therefore, we conclude that the Dirichlet-type space $D(\mu)$ coincides with de Branges-Rovnyak space $H(B)$ with equality of norms where $B=(b_{1},b_{2})$ and $b_{j}=\displaystyle\frac{p_{j}}{q}$, $j=1,2.$